OpenStudy (anonymous):

what are the solutions of 2x2 + 13x + 6 = 0.

OpenStudy (anonymous):

its 2x^2 by the way

OpenStudy (anonymous):

Common factoring: first, multiply the first and third term coefficients: That is, 2x6 =12 Now, find two numbers that multiply to 12 and add to 13: that is 12 and 1, because 12x1 = 12 and 12 + 1 =13. Then, as follows: 2x^2 + 12x + x + 6 Then, factor: 2x (x+6) + (x+6) (2x+1) (x+6) = 0 Now, you can find both of these values @ 0: 2x+1 = 0 2x = 1 x = 1/2 x+6 = 0 x = -6

OpenStudy (jiteshmeghwal9):

\[2x^2+12x+x+6=0\]\[2x(x+6)+(x+6)=0\]\[(2x+1)(x+6)=0\]so\[2x+1=0\]\[2x=-1\]\[x={-1 \over 2}\]&\[x+6=0\]\[x=-6\]

OpenStudy (anonymous):

Yeah, correct my first part: 2x+1=0, then 2x=-1, so it's -1/2, Sorry!

OpenStudy (anonymous):

OHHH its the factoring, i did wrong thanks so much guys

OpenStudy (anonymous):

OHHH its the factoring, i did wrong thanks so much guys

OpenStudy (anonymous):

2x^2+13x+6=0 break the middle term i.e 13x